OpenStudy (anonymous):

which inequality is equivalent to 6x-3y<=12 F:y<=2x-4 G:y>=2x+12 H:y>+=2x-4 J:y=-2x+4

OpenStudy (anonymous):

6x-3y<=12 -3y<=-6x+12 y>=2x-4 so you answer should be H